OpenStudy (anonymous):

26=m+5 solving a two step equation 6

OpenStudy (anonymous):

why exactly is there a floating 6? lol

OpenStudy (anonymous):

its like m over 6 

OpenStudy (anonymous):

sorry Im bad writing like that lol

OpenStudy (anonymous):

So, \(26 = \frac{ m }{ 6 } + 5\) ??

OpenStudy (anonymous):

yes

thomaster (thomaster):

Subtract 5 on both sides, then multiply everything by 6 to get rid of the fraction. That would give you the answer

OpenStudy (anonymous):

it gives me 3.5

thomaster (thomaster):

\(\large26 = \dfrac{ m }{ 6 } + 5\) subtract 5 on both sides \(\large21 = \dfrac{ m }{ 6 } \) now multiply both sides by 6. \(\dfrac{m}6*6 = m\) So \(\large m=21*6 \)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."